The area of a rectangle is 96 square centimeters. If the width of the rectangle is 4 centimeters less than its length, what are the dimensions of the rectangle?
Answer by ewatrrr(24785)

 
Hi,
width of the rectangle is 4 centimeters less than its length
Let x and (x-4) represent length and width respectively
Question states***
x(x-4)= 96 in^2
x^2 - 4x - 96 = 0
(x-12)(x+8) = 0
x = 12, throwing out negative solution for unit measure
Length is 12cm and width is 8cm
12in*8in = 96cm^2